# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ 7.1. Upper bound for $\text{IVM}^{\pm}[Q]$
The following example demonstrates our algorithm for $\text{IVM}^{\pm}[Q]$ from Theorem 3.4. Algorithm 1 in Appendix D.1 shows the general algorithm. Appendix D.2 shows how to extend it to $\text{IVM}^{\pm}_{\delta}[Q]$.
@@
Suppose we want to solve the problem $\text{IVM}^{\pm}[Q_{\triangle}]$ for $Q_{\triangle}$ in Eq. (1), where we have a stream of $N$ single-tuple inserts/deletes into $R,S$ and $T$. For simplicity, assume that $N$ is known in advance (Remark D.1 explains how to remove this assumption.) We introduce a new interval variable $[Z]$ and use it to represent time. Specifically, $[Z]$ will represent the lifespan of every tuple in the database, in the spirit of temporal databases (Jensen and Snodgrass, 2018). By adding $[Z]$ to every atom, we obtain the time extension $\overline{Q}_{\triangle}$ of $Q_{\triangle}$ in Eq. (11). Suppose that the $\tau$-th update is an insert of a tuple $(a,b)$ into $R$, i.e. $+R(a,b)$. Then, we apply the insert $+\overline{R}([\tau,\infty],a,b)$ into the time extension, indicating that the tuple $(a,b)$ lives in $R$ from time $\tau$ on (since we don’t know yet its future deletion time). Now suppose that the $\tau^{\prime}$-th update (for some $\tau^{\prime}>\tau$) is a delete of the same tuple $(a,b)$ in $R$. Then, we replace the tuple $([\tau,\infty],a,b)$ in $\overline{R}$ with $([\tau,\tau^{\prime}],a,b)$. We can extract the result of $Q_{\triangle}$ at any time $\tau$ by selecting output tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. Therefore, if we can efficiently maintain $\overline{Q}_{\triangle}$, we can efficiently maintain $Q_{\triangle}$. Table 2 shows the same stream of 8 updates into $Q_{\triangle}$ as Table 1 but adds the corresponding updates to $\overline{Q}_{\triangle}$. The final database $\overline{\mathcal{D}}$ after all updates have taken place is the same as the one shown in Figure 1(d).
@@ 8 $-R(a_{1},b_{1})$ $-\overline{R}([1,\infty],a_{1},b_{1})$, $+\overline{R}([1,8],a_{1},b_{1})$
$-\{(z_{1},z_{2},b_{1},c_{1})\mid(z_{1}\circ z_{2})\in\text{\sf CP}_{8}([2,8])\}$
Out of the box, Theorem 3.1 is only limited to inserts. However, we can show that these pairs of inserts/deletes have a special structure that allows us to maintain the guarantees of Theorem 3.1. In particular, at time 5, when we truncate a tuple $\bm{t}$ from $[2,8]$ to $[2,5]$, even after truncation, $\bm{t}$ still joins with the same set of tuples that it used to join with before the truncation. This is because the only intervals that overlap with $[2,8]$ but not with $[2,5]$ are the ones that start after 5. But these intervals are in the future, hence they don’t exist yet in the database (see Proposition D.2). We use this idea to amortize the cost of truncations over the inserts.
For $\text{IVM}^{\pm}[Q_{\triangle}]$, our target is to do constant-delay enumeration of the full output $Q_{\triangle}$ at time $\tau$. To that end, we enumerate tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. And to enumerate those, we enumerate tuples $(z_{1},z_{2},z_{3},a,b,c)$ from $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ where $z_{1}\circ z_{2}\circ z_{3}$ corresponds to an interval in $\text{\sf CP}_{N}([z])$ that contains $\tau$. This is a selection condition over $(z_{1},z_{2},z_{3})$. By construction, the tree decomposition of any component, say $\widehat{Q}_{123}$, must contain a bag $W$ that contains the variables $\{z_{1},z_{2},z_{3}\}$. This is because, by Definition 2.3, $\widehat{Q}_{123}$ contains an atom $\widehat{T}_{123}$ that contains these variables. We designate $W$ as the root of tree decomposition of $\widehat{Q}_{123}$ and start our constant-delay enumeration from $W$. The same goes for $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$. Moreover, the enumeration outputs of $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ are disjoint in this case. See the proof of Lemma D.3.
For $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$, our target is to enumerate the change to the output at time $\tau$. To that end, we enumerate tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ has $\tau$ as an endpoint. Appendix D.2 explains the technical challenges and how to address them.
@@ 7.2. Lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The following example demonstrates our lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$ from Theorem 3.6, which is based on a reduction from the static evaluation of a component $\widehat{Q}_{\bm{\sigma}}$ of the multivariate extension of $Q$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Algorithm 2 in Appendix D.4 details the reduction.
@@
Let us take an arbitrary database instance $\widehat{\mathcal{D}}_{123}=(\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123})$ and evaluate $\widehat{Q}_{123}$ over $\widehat{\mathcal{D}}_{123}$ using the oracle that solves $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$. Example C.6 shows how to reduce $\text{Eval}[\widehat{Q}_{123}]$ over $\widehat{\mathcal{D}}_{123}$ to the static evaluation of $\overline{Q}_{\triangle}$ from Eq. (11) over a database instance $\overline{\mathcal{D}}=(\overline{R},\overline{S},\overline{T})$ that satisfies:
@@
Moreover, $\overline{\mathcal{D}}$ can be constructed in linear time (Theorem C.9.) Next, we reduce the evaluation of $\overline{Q}_{\triangle}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$.
@@
Appendix D.4 highlights the challenges in extending the above lower bound from $\text{IVM}^{\pm}_{\delta}$ to $\text{IVM}^{\pm}$, which remains an open problem.
@@ 8. Conclusion
The assumptions that the queries have all variables free and the input database be initially empty can be lifted without difficulty. To support arbitrary conjunctive queries, our algorithms need to consider free-connex tree decompositions, which ensure constant-delay enumeration of the tuples in the query result. To support non-empty initial databases, we can use a preprocessing phase to construct our data structure, where all initial tuples get the same starting timestamp. The complexities stated in the paper need to be changed to account for the size of the initial database as part of the size of the update stream.

### v3

@@ 7.1. Upper bound for $\text{IVM}^{\pm}[Q]$
The following example demonstrates our algorithm for $\text{IVM}^{\pm}[Q]$ from Theorem 3.4. Algorithm 2 in Appendix E.1 shows the general algorithm. Appendix E.2 shows how to extend it to $\text{IVM}^{\pm}_{\delta}[Q]$.
@@
Suppose we want to solve the problem $\text{IVM}^{\pm}[Q_{\triangle}]$ for $Q_{\triangle}$ in Eq. (1), where we have a stream of $N$ single-tuple inserts/deletes into $R,S$ and $T$. For simplicity, assume that $N$ is known in advance (Remark E.1 explains how to remove this assumption.) We introduce a new interval variable $[Z]$ and use it to represent time. Specifically, $[Z]$ will represent the lifespan of every tuple in the database, in the spirit of temporal databases (Jensen and Snodgrass, 2018). By adding $[Z]$ to every atom, we obtain the time extension $\overline{Q}_{\triangle}$ of $Q_{\triangle}$ in Eq. (11). Suppose that the $\tau$-th update is an insert of a tuple $(a,b)$ into $R$, i.e. $+R(a,b)$. Then, we apply the insert $+\overline{R}([\tau,\infty],a,b)$ into the time extension, indicating that the tuple $(a,b)$ lives in $R$ from time $\tau$ on (since we don’t know yet its future deletion time). Now suppose that the $\tau^{\prime}$-th update (for some $\tau^{\prime}>\tau$) is a delete of the same tuple $(a,b)$ in $R$. Then, we replace the tuple $([\tau,\infty],a,b)$ in $\overline{R}$ with $([\tau,\tau^{\prime}],a,b)$. We can extract the result of $Q_{\triangle}$ at any time $\tau$ by selecting output tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. Therefore, if we can efficiently maintain $\overline{Q}_{\triangle}$, we can efficiently maintain $Q_{\triangle}$. Table 2 shows the same stream of 8 updates into $Q_{\triangle}$ as Table 1 but adds the corresponding updates to $\overline{Q}_{\triangle}$. The final database $\overline{\mathcal{D}}$ after all updates have taken place is the same as the one shown in Figure 2(d).
@@ 8 $-R(a_{1},b_{1})$ $-\overline{R}([1,\infty],a_{1},b_{1})$, $+\overline{R}([1,8],a_{1},b_{1})$
$-\{(z_{1},z_{2},b_{1},c_{1})\mid(z_{1}\circ z_{2})\in\text{\sf CP}_{8}([2,8])\}$ +$+\{(z_{1},z_{2},b_{1},c_{1})\mid(z_{1}\circ z_{2})\in\text{\sf CP}_{8}([2,5])\}$
Out of the box, Theorem 3.1 is only limited to inserts. However, we can show that these pairs of inserts/deletes have a special structure that allows us to maintain the guarantees of Theorem 3.1. In particular, at time 5, when we truncate a tuple $\bm{t}$ from $[2,8]$ to $[2,5]$, even after truncation, $\bm{t}$ still joins with the same set of tuples that it used to join with before the truncation. This is because the only intervals that overlap with $[2,8]$ but not with $[2,5]$ are the ones that start after 5. But these intervals are in the future, hence they don’t exist yet in the database (see Proposition E.2). We use this idea to amortize the cost of truncations over the inserts.
For $\text{IVM}^{\pm}[Q_{\triangle}]$, our target is to do constant-delay enumeration of the full output $Q_{\triangle}$ at time $\tau$. To that end, we enumerate tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. And to enumerate those, we enumerate tuples $(z_{1},z_{2},z_{3},a,b,c)$ from $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ where $z_{1}\circ z_{2}\circ z_{3}$ corresponds to an interval in $\text{\sf CP}_{N}([z])$ that contains $\tau$. This is a selection condition over $(z_{1},z_{2},z_{3})$. By construction, the tree decomposition of any component, say $\widehat{Q}_{123}$, must contain a bag $W$ that contains the variables $\{z_{1},z_{2},z_{3}\}$. This is because, by Definition 2.3, $\widehat{Q}_{123}$ contains an atom $\widehat{T}_{123}$ that contains these variables. We designate $W$ as the root of tree decomposition of $\widehat{Q}_{123}$ and start our constant-delay enumeration from $W$. The same goes for $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$. Moreover, the enumeration outputs of $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ are disjoint in this case. See the proof of Lemma E.3.
For $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$, our target is to enumerate the change to the output at time $\tau$. To that end, we enumerate tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ has $\tau$ as an endpoint. Appendix E.2 explains the technical challenges and how to address them.
@@ 7.2. Lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The following example demonstrates our lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$ from Theorem 3.6, which is based on a reduction from the static evaluation of a component $\widehat{Q}_{\bm{\sigma}}$ of the multivariate extension of $Q$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Algorithm 3 in Appendix E.4 details the reduction.
@@
Let us take an arbitrary database instance $\widehat{\mathcal{D}}_{123}=(\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123})$ and evaluate $\widehat{Q}_{123}$ over $\widehat{\mathcal{D}}_{123}$ using the oracle that solves $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$. Example D.6 shows how to reduce $\text{Eval}[\widehat{Q}_{123}]$ over $\widehat{\mathcal{D}}_{123}$ to the static evaluation of $\overline{Q}_{\triangle}$ from Eq. (11) over a database instance $\overline{\mathcal{D}}=(\overline{R},\overline{S},\overline{T})$ that satisfies:
@@
Moreover, $\overline{\mathcal{D}}$ can be constructed in linear time (Theorem D.9.) Next, we reduce the evaluation of $\overline{Q}_{\triangle}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$.
@@
Appendix E.4 highlights the challenges in extending the above lower bound from $\text{IVM}^{\pm}_{\delta}$ to $\text{IVM}^{\pm}$, which remains an open problem.
@@ 8. Conclusion
The assumptions that the queries have all variables free and the input database be initially empty can be lifted without difficulty. To support arbitrary conjunctive queries, our algorithms need to consider free-connex tree decompositions, which ensure constant-delay enumeration of the tuples in the query result. To support non-empty initial databases, we can use a preprocessing phase to construct our data structure, where all initial tuples get the same starting timestamp. The complexities stated in the paper need to be changed to account for the size of the initial database as part of the size of the update sequence.